The Navier-Stokes equation in distribution spaces

Anna L. Mazzucato Yale University and IMA

We discuss the Navier-Stokes equation in distribution spaces. We investigate existence and uniqueness by means of fixed-point methods and Littlewood-Paley theory. In particular, we consider the scale of Besov-Morrey spaces.